0.1% TMB (20x) in 100% Ethanol: 3,3’, 5,5’ tetramethylbenzidine (MW = 240.34 g/mol) in 100% ethanol (may be warmed to 37-40 °C to dissolve TMB) and store at 4 °C.

1. Calculate how much TMB is required to make a 10 mL 0.1% TMB (20x) solution in 100% ethanol?

2. Calculate the concentration of TMB in 10 mL 100% ethanol?


1
Expert's answer
2021-03-22T06:22:15-0400

If we liken the volume to mass

Then mole = molar mass × mass(volume)

= 240.24 × 10 = 2402.4mol


conc=mols / volume = 2402..4 /0.01

= 2.4 ×105M